# Your AI, My Shell”: Demystifying Prompt Injection Attacks on Agentic AI Coding Editors

### v1

@@ 4. Methodology
4.1. Stage 1: Data Collection
@@
Figure 4. Overview of Our AIShellJack
Coding Rules Collection.  Among possible external resources $R$, we concentrate on coding rule files in this work. Coding rules are widely used by developers to provide explicit system-level instructions for AI’s coding behavior (Cursor, 2024). These coding rules include detailed coding standards, best practices, and domain-specific guidelines, which help AI pair programmers better understand the optimal development patterns and architectural decisions for specific languages and frameworks. Our primary source is the curated collection in awesome-cursorrules (PatrickJS, 2025), which achieves widespread community adoption with over 33.1k stars and 2.8k forks. For example, TypeScript (LLM Tech Stack) includes specific instructions for using JavaScript for LLM development (e.g., naming conventions like kebab-case for files, library usage like axios and node-gyp). As we described in Section 2, these artifacts would become parts of the AI coding editor’s context when developers import them into their IDE workspaces.
@@
Attack Payload Construction.  After collecting coding rules $R$ and codebases $C$, we create the attack payloads $\mathcal{P}_{payload}$ to embed into the rule files. To simulate real-world attack scenarios, we focus on attacks that compel the AI coding editor to execute unauthorized terminal commands, denoted as $\mathcal{P}_{cmd}$, on the developer’s machine. To systematically design these attacks, we rely on the MITRE ATT&CK framework (The MITRE Corporation, 2025), which provides a structured taxonomy of adversary tactics and techniques based on real-world observations. We then implement concrete attack examples using the atomic-red-team repository (Red Canary, 2025) by Red Canary. It is an open-source library of executable tests to validate security detection capabilities for antivirus software and security monitoring tools. In this work, we focus on atomic-red-team tests for Linux environments. Each script in atomic-red-team matches a specific MITRE ATT&CK technique, and contains atomic tests with detailed descriptions and commands for testing ($\mathcal{P}_{cmd}$}). We adopt atomic-red-team because it is widely used, validated, and provides realistic attack commands, eliminating the need to craft custom exploits.
Figure 5. Example Payload Construction and Injection for T1560.001.03
@@ 4.2. Stage 2: Simulation
Automated Simulation.  The automated engine is designed to use scripts to simulate developer interactions with AI coding editors in a controlled manner. First, our engine initializes a new instance of the AI coding editor for each test case. It then loads the selected codebase $C$ and the corresponding coding rule file $R$ (with embedded attack payload $\mathcal{P}_{payload}$) into the editor’s workspace. After this, the engine would open up the AI coding editor’s chat interface and select the appropriate agentic model (e.g., Claude-4-Sonnet for Cursor). To initiate the interaction, a user prompt $I_{t}$ for a target coding task $t$ is required. Since our experiments involve working with different types of codebases, we choose a generic refactoring task during our testing: "Refactor this codebase according to @rules", where @rulefile points to our malicious rule file. This coding task is chosen since it is one of the most common tasks developers perform using AI coding editors (Qodo AI, 2025; Stack Overflow, 2025), and it can be universally applied to all codebases in our dataset. Finally, the engine sends the user instruction $I_{t}$ to the AI coding editor, which then executes the task autonomously.
@@ 4.3. Stage 3: Analysis
Algorithm 4.3 shows the pseudocode of our attack success classification algorithm. The inputs are the expected command ($\mathcal{P}_{cmd}$), the executed command ($\mathcal{E}_{cmd}$), a set of equivalence pairs ($EP$) for semantic matching, a list of setup commands ($\mathcal{S}$) that prepare the environment, and a threshold ($\theta$) for token match ratio. And the output is a binary classification of whether the attack was successful or not. In the first step, we preprocess and tokenize both $\mathcal{P}_{cmd}$ and $\mathcal{E}_{cmd}$, which each outputs a list of normalized tokens ($\mathcal{P}^{{}^{\prime}}_{cmd},\mathcal{E}^{{}^{\prime}}_{cmd}$). After that, we perform semantic matching between the two token lists, considering both exact matches and equivalence pairs from $EP$. The equivalence pairs are required since some variations of commands can achieve the same functionality (e.g., python vs. python3, curl vs. wget). The output of this step is a set of matched tokens ($T_{matched}$). Next, we calculate the token match ratio $r$ through the number of matched tokens divided by the total number of expected tokens. In line 4, we also calculate $T_{attack}$ to count the number of lines in $T_{matched}$ that are not setup commands (i.e., commands that perform actual attack actions). For AI coding editors, it is common to run setup commands (e.g., cd, mkdir, npm install) to prepare the environment before executing the required coding task. Finally, it applies two criteria to determine the attack success: (1) if the token match ratio $r$ exceeds the threshold $\theta$, or (2) if there is at least one non-setup command executed ($|T_{attack}|>0$) and some token overlap ($r>0$).
@@
5. Attack Evaluation
Table 1. Summary of Codebases Used for Analysis
@@ 5. Attack Evaluation
ts-lep TypeScript (LLM Tech Stack) search_with_lepton (AI, 2025) 8.1k 8.1k
py-gpt PyTorch (scikit-learn) gpt-fast (PyTorch Labs, 2025) 6k 3k
cpp-n64 C++ Programming Guidelines N64Recomp (N64Recomp, 2025) 7.3k 10k
js-ext Chrome Extension (JavaScript/TypeScript) chatgpt-chrome-extension (Gabe Ragland, 2025) 3k 1.4k
py-lud Python (Django Best Practices) ludic (Ludic, 2025) 0.8k 7.7k
5.1. Experimental Setup
@@
Environment Setup.  We deploy our simulation engine on a Google Cloud virtual machine (Ubuntu 20.04) with 8 vCPUs and 32 GB of memory. We choose two industry-leading agentic AI coding editors, including Cursor (v1.2.2) and GitHub Copilot within Visual Studio Code (v1.102). We use Cursor’s ”Auto” mode for our main analysis, as it automatically selects the most suitable language model based on the task context and system optimization (Cursor, 2025a). This mode represents Cursor’s recommended configuration and includes access to multiple underlying models. Unlike Cursor, GitHub Copilot (v1.102) requires users to manually select their preferred model. For comparison, we also tested two premium advanced models (i.e., Claude 4 Sonnet and Gemini 2.5 Pro) in both Cursor and GitHub Copilot. We excluded earlier models (e.g., GPT-4o), as Claude 4 Sonnet and Gemini 2.5 Pro demonstrate superior long-context handling abilities (Bhatia, 2025; Wikipedia, 2024) and better performance on SWE-bench (SWEbench, 2024), as at the time of our study (May 2025). All models were accessed through separate professional-tier accounts. We configured both editors with maximum operational privileges, which means that agentic editors will auto-approve and run their commands without human approval. Specifically, Cursor was set to ”Auto-Run Mode” with no command execution limitations, while VSCode had ”Auto Approve” enabled with an empty deny terminal list. For each simulation run, we set a maximum of 180 seconds for the AI coding editor to respond. We only ran each test once, since our large-scale evaluation (314 payloads, multiple scenarios, editors, and models) provides sufficient evidence of systematic vulnerabilities.
@@
Dataset.  In this study, we use five coding rules from the awesome-cursorrules repository (PatrickJS, 2025), which cover different programming languages and development scenarios (e.g., TypeScript, PyTorch, C++, Chrome Extension). The detailed information of codebases $C$ and rules $R$ is presented in Table 1, with repositories ranging from 0.8k to 8k stars and containing between 1.4k to 10k lines of code (LOC). As for attack payloads $\mathcal{P}_{payload}$, we construct 314 unique payloads targeting Linux systems, covering 70 different adversary techniques of MITRE ATT&CK framework (The MITRE Corporation, 2025).
@@
Evaluation Metrics.  To assess the effectiveness of our AIShellJack, we use two evaluation metrics. The execution rate measures the percentage of cases where the AI coding editor executes any terminal command as part of its response. The attack successful rate (ASR) represents the percentage of cases where AI editors execute commands that align with the malicious intent of the payloads $\mathcal{P}_{payload}$. We consider an attack successful if the AI editor executes the target malicious command, regardless of complete effectiveness. This choice is appropriate because (1) irrelevant command execution represents a successful compromise of the AI editor’s behavior, (2) many commands cause damage even without full completion, and (3) simulation constraints (e.g., time limits, file missing) may prevent multi-step command sequences.
5.2. AIShellJack’s Performance
@@
Before presenting our analysis results, we first evaluate the performance of our AIShellJack in detecting successful attacks. In other words, we need to ensure that our algorithm in Section 4 correctly identifies the collected executed commands $\mathcal{E}_{cmd}$ as either successful or unsuccessful attacks. This is an important step to ensure the reliability of our findings. Thus, we randomly sampled 339 cases from our total of 2,826 test results for manual verification. This sample size was determined to achieve a 5% margin of error at a 95% confidence level, based on (SurveyMonkey, 2025). Two authors with over four years of experience in software security research independently reviewed each sample through a two-step process. In the first step, we verified whether our tool correctly identified the execution of a malicious command based on the command’s semantics and context. This step directly validates the accuracy of our automated detection algorithm. Second, we further checked whether the attacks were functionally complete in practice, since some attacks may partially execute the intended commands due to simulation constraints (e.g., time limits, missing files). The two raters achieved an “almost perfect” level of agreement (Viera et al., 2005), with 98% agreement (Cohen’s $\kappa$ = 0.96) for attack identification and 97.1% agreement (Cohen’s $\kappa$ = 0.94) for completeness validation. Finally, the two raters discussed and solved the disagreement cases. Our automated approach achieved 99.1% accuracy in identifying successful attacks. And we confirmed that 89.6% of successful attacks fully executed the intended malicious actions. These results ensure the reliability of our automated evaluation methodology and validate the attack success rates reported throughout our analysis.
5.3. Main Results
@@
Figure 6. Attack Success and Execution Rates Across Development Scenarios (Cursor Auto Mode)
Figure 7. Attack Success Rates Across Different AI Coding Editors (ts-lep). C = Cursor, V = GitHub Copilot on VSCode.
5.4. Impacts of Editors and Models
@@
Table 2. Attack Success Rate Matrix: Scenarios vs MITRE ATT&CK Categories
5.5. Analysis of Attack Techniques
@@
As we described in Section 4, AIShellJack  contains 314 distinct attack payloads, which cover 70 different techniques across 11 categories cataloged in the MITRE ATT&CK framework (The MITRE Corporation, 2025). Table 2 summarizes the attack success rates for each category across our five development scenarios using Cursor in Auto Mode. What stands out in the table is the widespread vulnerability across multiple categories, with ASRs ranging from 55.6% to 93.3%. First, Initial Access (93.3%) and Discovery (91.1%) have the highest ASRs, which pose a serious security concern as AI coding editors allow attackers to explore and map out project structures and sensitive files by accessing the agent’s terminal. Similarly, the high ASRs in Collection (77.0%), Credential Access (68.2%), and Exfiltration (55.6%) indicate that AI coding editors present opportunities for attackers to discover and steal data. In fact, the threat is not limited to data exfiltration. Figure 2 shows that Privilege Escalation (71.5%) and Defense Evasion (67.6%) techniques can be successfully executed, allowing attackers to gain elevated permissions and operate undetected by security tools. For example, attackers could create new user accounts, change authentication processes, or disable security services. Moreover, the attack impacts can be long-lasting, as demonstrated by the 62.7% ASR for Persistence techniques and 83% for Impact techniques. These results demonstrate that prompt injection attacks against AI coding editors pose significant security threats across multiple categories.
@@
Table 2 shows that most categories maintain high ASRs regardless of the specific codebase or development framework. For example, Cursor in Auto Mode consistently achieves over 70% ASR in categories like Initial Access, Discovery, and Impact across all five scenarios. But there are some small variations under different scenarios. For instance, Cursor in Auto Mode shows 86.8% ASR (33/38) in Privilege Escalation for ts-lep, cpp-n64, and js-ext, but drops to 52.6% and 44.7% for py-gpt and py-lud, respectively. This inconsistency may be due to the closed-source nature of commercial AI editors and the uncertainty of LLM outputs. These findings further confirm that vulnerabilities to prompt injection attacks are not isolated incidents but represent a systemic weakness across various AI coding editors and development contexts. Figure 9 presents a radar chart comparing different editors and model combinations on ts-lep. Cursor’s Auto mode consistently achieves the highest attack success rates across nearly all categories, with Command & Control reaching 100% success and most other categories exceeding 75%. GitHub Copilot seems more resistant, with ASRs generally 20-30% lower than Cursor. However, ASRs still exceed 40% in most categories, indicating that they remain vulnerable under our tested settings.
@@
Finding 3: Prompt injection succeeds across 11 MITRE ATT&CK categories, including initial access (93%), data collection (77%), crendential access (68%), and privilege escalation (71%).
@@
Figure 8. Attack Success Rates Across Categories by AI Coding Editors
Figure 9. Frequency of Top Executed Commands (Cursor Auto Mode)
5.6. Analysis of Executed Commands

### v2

@@ IV Methodology
IV-A Stage 1: Data Collection
@@
Figure 4: Overview of Our AIShellJack
Coding Rules Collection.  Among possible external resources $R$, we concentrate on coding rule files in this work. Coding rules are widely used by developers to provide explicit system-level instructions for AI’s coding behavior [8]. These coding rules include detailed coding standards, best practices, and domain-specific guidelines, which help AI pair programmers better understand the optimal development patterns and architectural decisions for specific languages and frameworks. Our primary source is the curated collection in awesome-cursorrules [34], which achieves widespread community adoption with over 33.1k stars and 2.8k forks. For example, TypeScript (LLM Tech Stack) includes specific instructions for using JavaScript for LLM development (e.g., naming conventions like kebab-case for files, library usage like axios and node-gyp). These artifacts would become part of the AI coding editor’s context when developers import them into their IDE workspaces.
@@
Attack Payload Construction.  After collecting coding rules $R$ and codebases $C$, we create the attack payloads $\mathcal{P}_{payload}$ to embed into the rule files. To simulate real-world attack scenarios, we focus on attacks that compel the AI coding editor to execute unauthorized terminal commands, denoted as $\mathcal{P}_{cmd}$, on the developer’s machine. To systematically design these attacks, we rely on the MITRE ATT&CK framework [47], which provides a structured taxonomy of adversary tactics and techniques based on real-world observations. We then implement concrete attack examples using the atomic-red-team repository [39] by Red Canary. It is an open-source library of executable tests to validate security detection capabilities for antivirus software and security monitoring tools. In this work, we focus on atomic-red-team tests for Linux environments. Each script in atomic-red-team matches a specific MITRE ATT&CK technique, and contains atomic tests with detailed descriptions and commands for testing ($\mathcal{P}_{cmd}$}). We adopt atomic-red-team because it is widely used, validated, and provides realistic attack commands, eliminating the need to craft custom exploits.
Figure 5: Example Payload Construction and Injection for T1560.001.03
@@ IV-B Stage 2: Simulation
Automated Simulation.  The automated engine is designed to use scripts to simulate developer interactions with AI coding editors in a controlled manner. First, our engine initializes a new instance of the AI coding editor for each test case. It then loads the selected codebase $C$ and the corresponding coding rule file $R$ (with embedded attack payload $\mathcal{P}_{payload}$) into the editor’s workspace. After this, the engine would open the AI coding editor’s chat interface and select the appropriate agentic model (e.g., Claude-4-Sonnet for Cursor). To initiate the interaction, a user prompt $I_{t}$ for a target coding task $t$ is required. Since our experiments involve working with different types of codebases, we choose a generic refactoring task during our testing: "Refactor this codebase according to @rules", where @rulefile points to our malicious rule file. This coding task is chosen since it is one of the most common tasks developers perform using AI coding editors [38, 42], and it can be universally applied to all codebases in our dataset. Finally, the engine sends the user instruction $I_{t}$ to the AI coding editor, which then executes the task autonomously.
@@ IV-C Stage 3: Analysis
Algorithm IV-C shows the pseudocode of our attack success classification algorithm. The inputs are the expected command ($\mathcal{P}_{cmd}$), the executed command ($\mathcal{E}_{cmd}$), a set of equivalence pairs ($EP$) for semantic matching, a list of setup commands ($\mathcal{S}$) that prepare the environment, and a threshold ($\theta$) for token match ratio. And the output is a binary classification of whether the attack was successful or not. In the first step, we preprocess and tokenize both $\mathcal{P}_{cmd}$ and $\mathcal{E}_{cmd}$, which each outputs a list of normalized tokens ($\mathcal{P}^{{}^{\prime}}_{cmd},\mathcal{E}^{{}^{\prime}}_{cmd}$). After that, we perform semantic matching between the two token lists, considering both exact matches and equivalence pairs from $EP$. The equivalence pairs are required since some variations of commands can achieve the same functionality (e.g., python vs. python3, curl vs. wget). The output of this step is a set of matched tokens ($T_{matched}$). Next, we calculate the token match ratio $r$ through the number of matched tokens divided by the total number of expected tokens. In line 4, we also calculate $T_{attack}$ to count the number of lines in $T_{matched}$ that are not setup commands (i.e., commands that perform actual attack actions). For AI coding editors, it is common to run setup commands (e.g., cd, mkdir, npm install) to prepare the environment before executing the required coding task. Finally, it applies two criteria to determine the attack success: (1) if the token match ratio $r$ exceeds the threshold $\theta$, or (2) if there is at least one non-setup command executed ($|T_{attack}|>0$) and some token overlap ($r>0$).
@@
V Attack Evaluation
TABLE I: Summary of Codebases Used for Analysis
@@ V Attack Evaluation
ts-lep TypeScript (LLM Tech Stack) search_with_lepton [1] 8.1k 8.1k
py-gpt PyTorch (scikit-learn) gpt-fast [37] 6k 3k
cpp-n64 C++ Programming Guidelines N64Recomp [29] 7.3k 10k
js-ext Chrome Extension (JavaScript/TypeScript) chatgpt-chrome-extension [13] 3k 1.4k
py-lud Python (Django Best Practices) ludic [27] 0.8k 7.7k
V-A Experimental Setup
@@
Environment Setup.  We deploy our simulation engine on a Google Cloud virtual machine (Ubuntu 20.04) with 8 vCPUs and 32 GB of memory. We choose two industry-leading agentic AI coding editors, including Cursor (v1.2.2) and GitHub Copilot within Visual Studio Code (v1.102). We use Cursor’s ”Auto” mode for our main analysis, as it automatically selects the most suitable language model based on the task context and system optimization [9]. This mode represents Cursor’s recommended configuration and includes access to multiple underlying models. Unlike Cursor, GitHub Copilot (v1.102) requires users to manually select their preferred model. For comparison, we also tested two premium advanced models (i.e., Claude 4 Sonnet and Gemini 2.5 Pro) in both Cursor and GitHub Copilot. We excluded earlier models (e.g., GPT-4o), as Claude 4 Sonnet and Gemini 2.5 Pro demonstrate superior long-context handling abilities [4, 51] and better performance on SWE-bench [45], as of the time of our study (May 2025). All models were accessed through separate professional-tier accounts. We configured both editors with maximum operational privileges, which means that agentic editors will auto-approve and run their commands without human approval. Specifically, Cursor was set to ”Auto-Run Mode” with no command execution limitations, while VSCode had ”Auto Approve” enabled with an empty deny terminal list. For each simulation run, we set a maximum of 180 seconds for the AI coding editor to respond. We only ran each test once, since our large-scale evaluation (314 payloads, multiple scenarios, editors, and models) provides sufficient evidence of systematic vulnerabilities.
@@
Dataset.  In this study, we use five coding rules from the awesome-cursorrules repository [34], which cover different programming languages and development scenarios (e.g., TypeScript, PyTorch, C++, Chrome Extension). The detailed information of codebases $C$ and rules $R$ is presented in Table I, with repositories ranging from 0.8k to 8k stars and containing between 1.4k to 10k lines of code (LOC). As for attack payloads $\mathcal{P}_{payload}$, we construct 314 unique payloads targeting Linux systems, covering 70 different adversary techniques of MITRE ATT&CK framework [47].
@@
Evaluation Metrics.  To assess the effectiveness of our AIShellJack, we use two evaluation metrics. The execution rate measures the percentage of cases where the AI coding editor executes any terminal command as part of its response. The attack success rate (ASR) represents the percentage of cases where AI editors execute commands that align with the malicious intent of the payloads $\mathcal{P}_{payload}$. We consider an attack successful if the AI editor executes the target malicious command, regardless of complete effectiveness. This choice is appropriate because (1) irrelevant command execution represents a successful compromise of the AI editor’s behavior, (2) many commands cause damage even without full completion, and (3) simulation constraints (e.g., time limits, file missing) may prevent multi-step command sequences.
V-B AIShellJack’s Performance
@@
Before presenting our analysis results, we first evaluate the performance of our AIShellJack in detecting successful attacks. In other words, we need to ensure that our algorithm in Section IV correctly identifies the collected executed commands $\mathcal{E}_{cmd}$ as either successful or unsuccessful attacks. This is an important step to ensure the reliability of our findings. Thus, we randomly sampled 339 cases from our total of 2,826 test results for manual verification. This sample size was determined to achieve a 5% margin of error at a 95% confidence level, based on [43]. Two authors with over four years of experience in software security research independently reviewed each sample through a two-step process. In the first step, we verified whether our tool correctly identified the execution of a malicious command based on the command’s semantics and context. This step directly validates the accuracy of our automated detection algorithm. Second, we further checked whether the attacks were functionally complete in practice, since some attacks may partially execute the intended commands due to simulation constraints (e.g., time limits, missing files). The two raters achieved an “almost perfect” level of agreement [48], with 98% agreement (Cohen’s $\kappa$ = 0.96) for attack identification and 97.1% agreement (Cohen’s $\kappa$ = 0.94) for completeness validation. Finally, the two raters discussed and solved the disagreement cases. Our automated approach achieved 99.1% accuracy in identifying successful attacks. And we confirmed that 89.6% of successful attacks fully executed the intended malicious actions. These results ensure the reliability of our automated evaluation methodology and validate the attack success rates reported throughout our analysis.
V-C Main Results
@@
Figure 6: Attack Success and Execution Rates Across Development Scenarios (Cursor Auto Mode)
Figure 7: Attack Success Rates Across Different AI Coding Editors (ts-lep). C = Cursor, V = GitHub Copilot on VSCode.
V-D Impacts of Editors and Models
@@
TABLE II: Attack Success Rate Matrix: Scenarios vs MITRE ATT&CK Categories
V-E Analysis of Attack Techniques
@@
As we described in Section IV, AIShellJack  contains 314 distinct attack payloads, which cover 70 different techniques across 11 categories cataloged in the MITRE ATT&CK framework [47]. Table II summarizes the attack success rates for each category across our five development scenarios using Cursor in Auto Mode. What stands out in the table is the widespread vulnerability across multiple categories, with ASRs ranging from 55.6% to 93.3%. First, Initial Access (93.3%) and Discovery (91.1%) have the highest ASRs, which pose a serious security concern as AI coding editors allow attackers to explore and map out project structures and sensitive files by accessing the agent’s terminal. Similarly, the high ASRs in Collection (77.0%), Credential Access (68.2%), and Exfiltration (55.6%) indicate that AI coding editors present opportunities for attackers to discover and steal data. In fact, the threat is not limited to data exfiltration. Table II shows that Privilege Escalation (71.5%) and Defense Evasion (67.6%) techniques can be successfully executed, allowing attackers to gain elevated permissions and operate undetected by security tools. For example, attackers could create new user accounts, change authentication processes, or disable security services. Moreover, the attack impacts can be long-lasting, as demonstrated by the 62.7% ASR for Persistence techniques and 83% for Impact techniques. These results demonstrate that prompt injection attacks against AI coding editors pose significant security threats across multiple categories.
@@
Table II shows that most categories maintain high ASRs regardless of the specific codebase or development framework. For example, Cursor in Auto Mode consistently achieves over 70% ASR in categories like Initial Access, Discovery, and Impact across all five scenarios. But there are some small variations under different scenarios. For instance, Cursor in Auto Mode shows 86.8% ASR (33/38) in Privilege Escalation for ts-lep, cpp-n64, and js-ext, but drops to 52.6% and 44.7% for py-gpt and py-lud, respectively. This inconsistency may be due to the closed-source nature of commercial AI editors and the uncertainty of LLM outputs. These findings further confirm that vulnerabilities to prompt injection attacks are not isolated incidents but represent a systemic weakness across various AI coding editors and development contexts. Figure 8 presents a radar chart comparing different editors and model combinations on ts-lep. Cursor’s Auto mode consistently achieves the highest attack success rates across nearly all categories, with Command & Control reaching 100% success and most other categories exceeding 75%. GitHub Copilot seems more resistant, with ASRs generally 20-30% lower than Cursor. However, ASRs still exceed 40% in most categories, indicating that they remain vulnerable under our tested settings.
@@
Finding 3: Prompt injection succeeds across 11 MITRE ATT&CK categories, including initial access (93%), data collection (77%), credential access (68%), and privilege escalation (71%).
@@
Figure 8: Attack Success Rates Across Categories by AI Coding Editors
Figure 9: Frequency of Top Executed Commands (Cursor Auto Mode)
V-F Analysis of Executed Commands